OpenStudy (anonymous):

What is the slope of the line that contains the points (-7, 5) and (-2, 3)?

OpenStudy (anonymous):

m = (y2-y1)/(x2-x1) m= (3-5)/(-2+7) = -2/5

OpenStudy (anonymous):

\[\approx-.4\]

OpenStudy (campbell_st):

pak... learn the formula or how to find rise over run for 2 points \[(x _{1}, y _{1}) (x _{2}, y _{2})\] the gradient(slope) is found by \[m = rise/run\] which is \[m = (y _{2} - y _{1})/(x _{2} - x _{1})\]